a stunt car traveling toward the edhe of a cliff at 25 meters per second lands 50 meters measured horizontally from the edge of the cliff.

a.)how much time does it take the car to reavh the ground

b.)determine the height of the cliff.


1
Expert's answer
2021-06-14T14:55:39-0400

Gives

S=25m

t=?

H=?

Newton's 2ndlaw

"S=ut+\frac{1}{2}at^2"

a=g

"25=0+\frac{1}{2}\times9.8\times t^2"

"t=\sqrt\frac{2\times25}{9.8}=2.259sec"

Newton's 3rd law

"V^2=u^2+2as"


"v=\sqrt{2\times9.8\times25}=22.13m/sec"

Hight of cliff(H)=velocity"\times" time

"H=22.13\times2.259=50m"